# MemFine: Memory-Aware Fine-Grained Scheduling for MoE Training

### v2

@@ 1 Introduction
With the exponential growth of artificial intelligence model sizes, traditional dense models face severe challenges in computation and memory wang2025step . The MoE model, as a sparse architecture, introduces a gating network and multiple expert sub-networks, significantly reducing the activated computation per forward pass while maintaining a massive number of model parameters shazeer2017outrageously . This characteristic makes it possible to build and train giant models with trillions or even tens of trillions of parameters under limited computational resources, rapidly becoming a key technical path driving the development of large language models and tasks in other domains zhang2022data ; yin2022deep ; zhang2023dual . Therefore, achieving efficient and stable large-scale MoE training has become an urgent demand in both industry and academia yang2025qwen3 ; rajbhandari2022deepspeed .
Although MoE models possess significant theoretical advantages, they face the challenge of load imbalance during practical training, especially in large-scale distributed environments cai2025survey . To maximize the expressive power and performance of the models, modern large-scale MoE training allows the number of tokens processed by each expert to be unrestricted. However, this strategy, combined with the stochastic routing of the gating network, leads to a highly non-uniform distribution of tokens among different experts lepikhin2020gshard ; wang2024auxiliary . This load imbalance phenomenon further translates into severe imbalance in GPU memory consumption. Specifically, during training, some GPUs might be assigned a number of tokens far exceeding their memory capacity, causing them to encounter memory overflow first and thus interrupting the entire training job. This issue severely constrains the potential for scaling MoE model size on GPU clusters with limited memory capacity kim2024scaling .
To address the aforementioned memory bottleneck, existing research primarily employs load-balancing solutions by setting a capacity factor to limit the number of tokens processed by each expert to a balanced level lepikhin2020gshard ; wang2024auxiliary ; fedus2022switch ; liu2024deepseek , while applying activation recomputation chen2016training ; korthikanti2023reducing and offloading strategies ren2021zero . Although this method can effectively avoid memory peaks, it severely disrupts the inherent dynamic routing mechanism of the MoE model, leading to degraded training accuracy and failing to realize the full potential of the MoE architecture. More importantly, these methods are generally applied on GPUs with large memory capacity and cannot effectively tackle the memory overflow problem caused by extreme load imbalance on GPUs with small memory capacity.
@@ 2.1 MoE Training and Expert Parallelism
The MoE model has emerged as a key technology in recent years to address the scalability challenges of ultra-large-scale models cai2025survey . Its core idea involves replacing the traditional dense feedforward network layer with a structure composed of multiple experts, i.e., small feedforward networks, and introducing a gating network that dynamically selects and activates a subset of these experts for each input token. This conditional computation paradigm enables a dramatic increase in the total number of model parameters while keeping the actual computational cost per token constant, thereby achieving an effective trade-off between computational efficiency and model capacity. Representative works yang2025qwen3 ; lepikhin2020gshard ; wang2024auxiliary ; fedus2022switch ; liu2024deepseek have successfully scaled the model parameter count to the trillion-level.
To efficiently train and deploy MoE models, EP has been developed as a specialized model parallelism strategy jin2025megascale . Unlike traditional model parallelism, which splits a single model layer across different devices, EP distributes different experts to distinct computing devices. The routing decisions of the gating network determine the flow of input data across devices, with each device only computing the experts it hosts. This paradigm has been adopted and optimized by mainstream open-source frameworks rasley2020deepspeed ; shoeybi2019megatron ; zheng2024llamafactory .
@@ 2.2 Load Unbalance
Current mainstream approaches primarily involve introducing an auxiliary load balancing loss into the loss function to regularize the gating network. The core idea of these methods is to encourage the average activation frequency of all experts to become more uniform. A classic early work is the auxiliary load balancing loss adopted in Switch Transformer fedus2022switch . It calculates the within-batch average routing probability for each expert and compares it with the average probability across all experts, using a mean square error loss to penalize deviations, thereby smoothly steering routing towards less loaded experts. To mitigate the accuracy impact of the auxiliary loss, DeepSeek wang2024auxiliary ; liu2024deepseek proposed an auxiliary loss-free load balancing algorithm, which dynamically adjusts the routing bias of experts to control token distribution. While these methods can effectively alleviate the overall load imbalance phenomenon, they still cannot avoid a small number of iterations with extreme load imbalance.
Another important approach combines hard constraints with soft incentives. For example, GShard lepikhin2020gshard , alongside an auxiliary loss, introduced the concept of expert capacity, setting an upper limit on the number of tokens each expert can process. When an expert reaches its capacity limit, excess tokens are forcibly routed to less loaded experts. This serves as an effective hard guarantee mechanism to prevent individual expert overload, but this method has been verified to affect the convergence of the overall training loss tang2025pangu .
Furthermore, activation memory optimization techniques such as full recomputation and activation offloading are often combined with the aforementioned methods to save memory and reduce the impact of distribution imbalance liu2024deepseek .
@@ 12 $2D_{t}bs^{\prime}g_{e}/(tc)$
This section introduces the theoretical memory cost model for MoE to elucidate the memory usage trends during MoE training. Generally, the GPU memory occupied during MoE training can be divided into two components zhao2025mofa :
@@ 5.1 Experimental Setup
  1. Method 1: Training without chunk splitting and applying full recomputation to save memory, which is widely used in Megatron-LM shoeybi2019megatron ;
@@ 7 Authors
Shanghai INESA(Group) Co., Ltd.:
Yue Sun, Hongxing Niu, Zhaoqun Sun, Hui Shen.

### v4

@@ 1 Introduction
With the exponential growth of artificial intelligence model sizes, traditional dense models face severe challenges in computation and memory Wang et al. (2025). The MoE model, as a sparse architecture, introduces a gating network and multiple expert sub-networks, significantly reducing the activated computation per forward pass while maintaining a massive number of model parameters Shazeer et al. (2017). This characteristic makes it possible to build and train giant models with trillions or even tens of trillions of parameters under limited computational resources, rapidly becoming a key technical path driving the development of large language models and tasks in other domains Zhang et al. (2022); Yin et al. (2022); Zhang et al. (2023). Therefore, achieving efficient and stable large-scale MoE training has become an urgent demand in both industry and academia Yang et al. (2025); Rajbhandari et al. (2022).
Although MoE models possess significant theoretical advantages, they face the challenge of load imbalance during practical training, especially in large-scale distributed environments Cai et al. (2025). To maximize the expressive power and performance of the models, modern large-scale MoE training allows the number of tokens processed by each expert to be unrestricted. However, this strategy, combined with the stochastic routing of the gating network, leads to a highly non-uniform distribution of tokens among different experts Lepikhin et al. (2020); Wang et al. (2024). This load imbalance phenomenon further translates into severe imbalance in GPU memory consumption. Specifically, during training, some GPUs might be assigned a number of tokens far exceeding their memory capacity, causing them to encounter memory overflow first and thus interrupting the entire training job. This issue severely constrains the potential for scaling MoE model size on GPU clusters with limited memory capacity Kim et al. (2024).
To address the aforementioned memory bottleneck, existing research primarily employs load-balancing solutions by setting a capacity factor to limit the number of tokens processed by each expert to a balanced level Lepikhin et al. (2020); Wang et al. (2024); Fedus et al. (2022); Liu et al. (2024), while applying activation recomputation Chen et al. (2016); Korthikanti et al. (2023) and offloading strategies Ren et al. (2021). Although this method can effectively avoid memory peaks, it severely disrupts the inherent dynamic routing mechanism of the MoE model, leading to degraded training accuracy and failing to realize the full potential of the MoE architecture. More importantly, these methods are generally applied on GPUs with large memory capacity and cannot effectively tackle the memory overflow problem caused by extreme load imbalance on GPUs with small memory capacity.
@@ 2.1 MoE Training and Expert Parallelism
The MoE model has emerged as a key technology in recent years to address the scalability challenges of ultra-large-scale models Cai et al. (2025). Its core idea involves replacing the traditional dense feedforward network layer with a structure composed of multiple experts, i.e., small feedforward networks, and introducing a gating network that dynamically selects and activates a subset of these experts for each input token. This conditional computation paradigm enables a dramatic increase in the total number of model parameters while keeping the actual computational cost per token constant, thereby achieving an effective trade-off between computational efficiency and model capacity. Representative works Yang et al. (2025); Lepikhin et al. (2020); Wang et al. (2024); Fedus et al. (2022); Liu et al. (2024) have successfully scaled the model parameter count to the trillion-level.
To efficiently train and deploy MoE models, EP has been developed as a specialized model parallelism strategy Jin et al. (2025). Unlike traditional model parallelism, which splits a single model layer across different devices, EP distributes different experts to distinct computing devices. The routing decisions of the gating network determine the flow of input data across devices, with each device only computing the experts it hosts. This paradigm has been adopted and optimized by mainstream open-source frameworks Rasley et al. (2020); Shoeybi et al. (2019); Zheng et al. (2024).
@@ 2.2 Load Unbalance
Current mainstream approaches primarily involve introducing an auxiliary load balancing loss into the loss function to regularize the gating network. The core idea of these methods is to encourage the average activation frequency of all experts to become more uniform. A classic early work is the auxiliary load balancing loss adopted in Switch Transformer Fedus et al. (2022). It calculates the within-batch average routing probability for each expert and compares it with the average probability across all experts, using a mean square error loss to penalize deviations, thereby smoothly steering routing towards less loaded experts. To mitigate the accuracy impact of the auxiliary loss, DeepSeek Wang et al. (2024); Liu et al. (2024) proposed an auxiliary loss-free load balancing algorithm, which dynamically adjusts the routing bias of experts to control token distribution. While these methods can effectively alleviate the overall load imbalance phenomenon, they still cannot avoid a small number of iterations with extreme load imbalance.
Another important approach combines hard constraints with soft incentives. For example, GShard Lepikhin et al. (2020), alongside an auxiliary loss, introduced the concept of expert capacity, setting an upper limit on the number of tokens each expert can process. When an expert reaches its capacity limit, excess tokens are forcibly routed to less loaded experts. This serves as an effective hard guarantee mechanism to prevent individual expert overload, but this method has been verified to affect the convergence of the overall training loss Tang et al. (2025).
Furthermore, activation memory optimization techniques such as full recomputation and activation offloading are often combined with the aforementioned methods to save memory and reduce the impact of distribution imbalance Liu et al. (2024).
@@ 12 $2D_{t}bs^{\prime}g_{e}/(tc)$
This section introduces the theoretical memory cost model for MoE to elucidate the memory usage trends during MoE training. Generally, the GPU memory occupied during MoE training can be divided into two components Zhao et al. (2025):
@@ 5.1 Experimental Setup
  1. Method 1: Training without chunk splitting and applying full recomputation to save memory, which is widely used in Megatron-LM Shoeybi et al. (2019);
@@ 7 Authors
INESA-ZTE Joint Laboratory:
Hui Shen, Zongtai Luo, Zhaoqun Sun, Hongxing Niu, Yue Sun.
Acknowledgement
We wish to express our sincere gratitude to the INESA-ZTE Joint Laboratory, which kindly provided the essential test environment and valuable technical assistance throughout the research and manuscript preparation process.